OpenStudy (anonymous):

The function f(x)=(3x−3)e−6x has one critical number. Find it.

jimthompson5910 (jim_thompson5910):

is the function this? \[\Large f(x) = (3x-3)e^{-6x}\]

OpenStudy (anonymous):

ya

jimthompson5910 (jim_thompson5910):

Use the product rule to get \[\Large f(x) = (3x-3)e^{-6x}\] \[\Large \frac{d}{dx}[f(x)] = \frac{d}{dx}[(3x-3)e^{-6x}]\] \[\Large f \ '(x) = \frac{d}{dx}[(3x-3)]e^{-6x}+(3x-3)\frac{d}{dx}[e^{-6x}]\] \[\Large f \ '(x) = 3e^{-6x}+(3x-3)*(-6e^{-6x})\] \[\Large f \ '(x) = 3e^{-6x}-18xe^{-6x}+18e^{-6x}\] \[\Large f \ '(x) = -18xe^{-6x}+21e^{-6x}\] \[\Large f \ '(x) = -3e^{-6x}(6x-7)\]

jimthompson5910 (jim_thompson5910):

Now plug in f ' (x) = 0 and solve for x \[\Large f \ '(x) = -3e^{-6x}(6x-7)\] \[\Large 0 = -3e^{-6x}(6x-7)\] I'll let you take over from here

OpenStudy (anonymous):

still not getting it

jimthompson5910 (jim_thompson5910):

Do you agree with how I got \(\Large f \ '(x)\) ?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

and you agree that we set f ' (x) equal to zero? or no?

jimthompson5910 (jim_thompson5910):

\[\Large f \ '(x) = -3e^{-6x}(6x-7)\] \[\Large 0 = -3e^{-6x}(6x-7)\] \[\Large -3e^{-6x}(6x-7) = 0\] \[\Large -3e^{-6x} = 0 \ \text{ or } \ 6x-7 = 0\] The equation \[\Large -3e^{-6x} = 0\] has no solutions. The equation \[\Large 6x-7 = 0\] has the solution x = _______ (fill in the blank)

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

np